OpenStudy (anonymous):

Solve the function 3x+2y=72 2x+4y=96

OpenStudy (luigi0210):

How would you like to tackle this problem?

OpenStudy (anonymous):

Um I don't know

OpenStudy (pablok777):

3x + 2y = 72 then multiply both members by 2 -6x - 4y = -144 and 2x + 4y = 96 add both equations -4x = -48 then x = 12 reemplazing int the first equation the x value 3*12 + 2*y = 72 then y = (72 - 36)/2 = 18

hero (hero):

I'd reduce that second equation a bit before starting.

OpenStudy (anonymous):

yes you can take '2' common from both sides

hero (hero):

So I would start with 3x + 2y = 72 x + 2y = 48

hero (hero):

Afterwards I would use elimination method to eliminate the y variable.

OpenStudy (luigi0210):

btw @pablok777 No doing others work for them or giving away answers